During the movement of an elevator , a screw in the overhead grill comes loose and falls to the floor of the elevator , which is 2.9 m below the grill. Compute the times of fall for the screw when the elevator is a). accelerating upward at a rate of 1/7 g . b). accelerating downwards at a rate of 1/7 g. c).moving at a constant speed.

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
According to the equation of motion,
S=ut+at2/2

u=0 here since it will fall from rest

a) S=ut+at2/2
    => 2.9=8gt2/14
   =>t=0.25s

b) S=ut+at2/2
    =>2.9=6gt2/7
   =>t=0.58s

c) S=ut+at2/2
   =>2.9=gt2/2
  =>t=0.762s


 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)